a speedboat travels 5 miles at 30 miles per hour, then 1 hour at 48 miles per hour. what was the boats average speed for the 6 mile trip?
Found 2 solutions by mananth, checkley77:
Answer by mananth(16949)

5 miles at 30 mph
distance traveled at this speed in 5 hour = 5*30 = 150 miles
..
1 hour at 48 mph.
distance traveled at this speed in 1 hour = 1*48 = 48 miles.
..
198 miles it travels in 6 hours
In 1 hour it travels 198/6 = 33 mph
..

Answer by checkley77(12844)

D=RT
5=30T
T=5/30
T=1/6 IS THE TIME FOR THE 5 MILE TRIP.
1+1/6=1.1667 IS THE TOTAL TIME.
5+48=53 IS THE TOTAL DISTANCE
53=R*1.1667
R=53/1.1667
R=45.43 MPH IS THE AVERAGE SPEED.